OpenStudy (anonymous):

Which is a counterexample that disproves the conjecture? If n is a positive integer, then 2n – 1 is a prime number. A. n = 6 B. n = 5 C. n = 3 D. n = 2 b

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@HelpBlahBlahBlah

ganeshie8 (ganeshie8):

Correct ! n = 5 : 2*5 - 1 = 10 - 1 = 9 9 is not a prime number